# Trade-off of Security, Latency, and Throughput of the Nakamoto Consensus

### v3

@@ Appendix A Proof of Theorem 1
Figure 9. Illustration of mining times of pacers and A-blocks after time $s$.
Under a confirmation rule of latency $t$, assume a target transaction is public by time $s$. Within this proof, we count block arrivals from time $s$, i.e., the first A-block is the first A-block mined after $s$, and the first pacer is the first pacer mined after $s$. Let $X_{0}$ denote the time from $s$ to the mining time of the first A-block. For $i=1,2,\dots$, let $X_{i}$ denote the inter-arrival time between the $i$-th A-block and the $(i+1)$-st A-block. Then, $X_{0},X_{1},\dots$ are independent and identically distributed (i.i.d.) exponential random variables with mean $\frac{1}{a}$. Let $Y_{0}$ denote the time from $s$ to the mining time of the first pacer. For $i=1,2,\dots$, let $Y_{i}$ denote the inter-arrival time between the $i$-th pacer and the $(i+1)$-st pacer. It is easy to see that $Y_{1},Y_{2},\dots$ are i.i.d. random variables with mean $\frac{1}{h}+\Delta$, and $Y_{0}$ has mean less than $\frac{1}{h}+\Delta$. An illustration of these random variables is given in Fig. 9.
@@
(45) $P(E_{j})=P(\sum_{i=0}^{j}X_{i}<\sum_{i=0}^{j}Y_{i},\sum_{i=0}^{j}X_{i}>t)$
(46) $\geq 1-P(\sum_{i=0}^{j}(X_{i}-Y_{i})\geq 0)-P(\sum_{i=0}^{j}X_{i}\leq t)$
$\geq 1-P(\sum_{i=1}^{j}(X_{i}-Y_{i})\geq-n)-P(X_{0}-Y_{0}\geq n)$
(47) $\quad-P(\sum_{i=0}^{j}X_{i}\leq t)$
@@
(48) $P(\sum_{i=1}^{j}(X_{i}-Y_{i})\geq-n)<\frac{\epsilon}{3}.$
@@ Appendix B Proof of Theorem 2
B.1. Critical Events
Figure 10. Illustration of all possible state transitions over time.
As a convention, we refer to the genesis block as the 0-th balancing candidate. We describe several types of events as follows: For every $i\in\{0,1,\cdots\}$, let
@@ B.1. Critical Events
If $U_{k}$ occurs, then at the end of the $k$-th candidacy, the candidate is not balanced, so the candidate becomes public and is strictly higher than all disagreeing blocks. As such, all honest mining power will extend chains that include the target transaction and are higher than the $k$-th candidate. This implies that the adversary must extend a shorter chain to catch up with the highest public honest block in order for the next candidate to be mined. In Appendix B.2, we show that the chance that no additional candidate will ever be mined is lower bounded as
(51) $P(A_{k}|U_{k})\geq\delta$
@@ B.2. Proof of (51)
(52) $P\left(\cap_{t\geq 0}\{A_{r,r+t}\leq P_{\Delta,t}\}\right)=P\left(\cap_{t\geq 0}\{A_{0,t}\leq P_{0,t}\}\right)$
(53) $=P\left(\max_{t\geq 0}\{A_{0,t}-P_{0,t}\}=0\right)$
(54) $=P(M=0).$
@@
(55) $P(A_{k}|U_{k})$ $\geq e^{-a\Delta}P(M=0)$
(56) $=\delta.$
B.3. Partially Markov Processes
@@
(60) $P^{m}(\{0\,b_{1}\dots b_{k-1}0\})=P(\{0\,b_{1}\dots b_{k-1}0\})$
for $k=1,\dots,m$ and as
@@
for $k=m+1,m+2,\dots$. The outcome $00\,b_{1}\dots b_{k-1}$ 0 with $k>m$ implies that $C_{m}$ occurs. It is straightforward to verify that (60), (61) defines a pmf. Evidently, $P^{m}$ is partially Markov in the sense that it is consistent with $P$ up until stage $m-1$, after which $P^{m}$ becomes Markov with the same transition probabilities as $P^{0}$. In particular, conditioned on $C_{m}$, the balanced stages before stage $m$ and the balanced stages afterward are independent.
For every pair of natural numbers $k$ and $j$, we define the following pair of random variables as functions of the sample path $00\,b_{1}\dots b_{k-1}$ 0:
(62) $L_{j}$ $=b_{1}+\dots+b_{\min(k,j)-1}$
(63) $R_{j}$ $=b_{j}+\dots+b_{k-1},$
@@
In Appendices B.4 and B.5, we show that for every $m=0,1,\dots$ and every $n=0,1,\dots$,
(65) $P^{m+1}(X\leq n)\geq P^{m}(X\leq n)$
@@
(70) $+P^{0}(R_{2}\geq n-1|B_{1})P^{0}(B_{1}|C_{1})$
(71) $=$ $(1-\delta)(1-\epsilon)f_{n}+\epsilon f_{n-1}.$
@@
(72) $f_{n}=\left(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\right)^{n}.$
@@
With confirmation by depth, there can be at most $k$ balanced heights before violation. In the remainder of this section, we are concerned with the number of balanced candidates up to the $k$-th candidacy, denoted as $X^{(k)}$. Conditioning on whether $C_{k+1}$ is visited, we can lower bound the cdf of $X^{(k)}$:
$P(X^{(k)}\leq n)$
(75) $\geq P^{0}(X^{(k)}\leq n)$
(76) $=P^{0}(X^{(k)}\leq n,X\geq n+1)+P^{0}(X\leq n)$
(77) $=P^{0}(X^{(k)}\leq n,X\geq n+1,C_{k+1})+P^{0}(X\leq n)$
where (75) can be easily seen from (65), and (77) holds because $X^{(k)}\leq n$ and $X\geq n+1$ implies the occurrence of $C_{k+1}$. Since we already know $P^{0}(X\leq n)$ by (74), what left is to calculate $P^{0}(X^{(k)}\leq n,X\geq n+1,C_{k+1})$. By law of total probability, we have
(78) $\begin{split}&P^{0}(X^{(k)}\leq n,X\geq n+1,C_{k+1})\\ &=\sum_{i=0}^{n}P^{0}(X^{(k)}=i,C_{k+1})P^{0}(X-X^{(k)}\geq n+1-i|C_{k+1}).\end{split}$
@@
(79) $=\sum_{i=0}^{n}\binom{k}{i}\epsilon^{i}((1-\epsilon)(1-\delta))^{k-i}(1-\delta)(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta})^{n-i+1}$
(80) $=(1-\delta)(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta})^{n-k+1}\sum_{i=0}^{n}\binom{k}{i}\epsilon^{i}(\frac{\epsilon(1-\epsilon)(1-\delta)}{\epsilon+\delta-\epsilon\delta})$
$=(1-\delta)(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta})^{n-k+1}(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta})^{k}$
(81) $\qquad\sum_{i=0}^{n}\binom{k}{i}(\epsilon+\delta-\epsilon\delta)^{i}(1-\epsilon)(1-\delta)^{k-i}$
(82) $=(1-\delta)(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta})^{n+1}F_{b}(n;k,(\epsilon+\delta-\epsilon\delta)),$
@@ B.4. Proof of (65)
(92) $\begin{split}&P^{m+1}(X\leq a)\\ &=P(L_{m}\leq a,C_{m}^{c})+\sum^{a}_{i=0}P(C_{m,i})P^{m+1}(X\leq a|C_{m,i}).\end{split}$
@@
(100) $\begin{split}P^{m}&(X\leq a|C_{m,i})=P^{0}(B_{m},C_{m+1}|C_{m,i})x_{1}\\ &+P^{0}(U_{m},C_{m+1}|C_{m,i})x_{2}+P^{0}(A_{m}|C_{m,i}).\end{split}$
@@
(103) $\lambda_{1}x_{1}+\lambda_{2}x_{2}+\lambda_{3}\leq\lambda_{1}^{\prime}x_{1}+\lambda_{2}^{\prime}x_{2}+\lambda_{3}^{\prime}.$
@@
$\lambda_{1}x_{1}$ $+\lambda_{2}x_{2}+\lambda_{3}$
(104) $=x_{2}+\lambda_{1}(x_{1}-x_{2})+\lambda_{3}(1-x_{2})$
(105) $\leq x_{2}+\lambda_{1}^{\prime}(x_{1}-x_{2})+\lambda_{3}^{\prime}(1-x_{2})$
(106) $=\lambda_{1}^{\prime}x_{1}+\lambda_{2}^{\prime}x_{2}+\lambda_{3}^{\prime}.$
@@
B.5. Proof of (66)
By the definition of limit, proving (66) equals to prove $\forall\lambda\geq 0,a\in\{0,1,\dots\},\exists N_{0},s.t.\forall N>N_{0}$, $|P^{N}(X\leq a)-P(X\leq a)|<\lambda$.
@@ B.5. Proof of (66)
$P^{N}(X\leq a)$ $=P^{N}(C_{N}^{c})P^{N}(X\leq a|C_{N}^{c})$
(109) $\quad+P^{N}(C_{N})P^{N}(X\leq a|C_{N}),$
(110) $P(X\leq a)$ $=P(C_{N}^{c})P(X\leq a|C_{N}^{c})+P(C_{N})P(X\leq a|C_{N}).$
By definition of (60) and (61), $P(X\leq a|C_{N}^{c})=P^{N}(X\leq a|C_{N}^{c})$ and $P(C_{N}^{c})=P^{N}(C_{N}^{c})$. Let $p_{1}=P^{N}(X\leq a|C_{N})$ and $p_{2}=P(X\leq a|C_{N}))$. We have,
$|P^{N}$ $(X\leq a)-P(X\leq a)|$
(111) $=P(C_{N})|p_{1}-p_{2}|$
(112) $\leq(1-\delta)\left(1-\delta(1-\epsilon)\right)^{N-1}|p_{1}-p_{2}|$
(113) $<\lambda,$
@@
(117) $P^{0}(C_{N})$ $=P^{0}(C_{1}|U_{0})\prod_{i=1}^{N-1}P^{0}(C_{i+1}|C_{i})$
(118) $=(1-\delta)\left(1-\delta(1-\epsilon)\right)^{N-1},$
@@ Appendix C Proof of Lemma 3
(127) $X_{0}=F_{X}^{-1}(U+F_{X}(-a)-\lfloor U+F_{X}(-a)\rfloor)$
(128) $Y_{0}=F_{Y}^{-1}(F_{Y}(a-1)-U-\lfloor F_{Y}(a-1)-U\rfloor)$
@@
(132) $F_{X}(b)$ $\geq F_{X}\circ X_{0}(u)$
(133) $=F_{X}\circ F^{-1}_{X}(u+F_{X}(-a))$
(134) $>u+F_{X}(-a)\quad a.e.,$
@@
(135) $F_{Y}(-b-1)$ $\geq F_{Y}\circ Y_{0}(u)$
(136) $=F_{Y}\circ F^{-1}_{Y}(F_{Y}(a-1)-u)$
(137) $>F_{Y}(a-1)-u\quad a.e.$

### v5

@@ VII Conclusion
Acknowledgement
D. Guo thanks Dr. David Tse for stimulating discussions that sparked the conceptualization of some initial ideas in the paper. This work was supported in part by the National Science Foundation (grant No. 2132700).
@@ Appendix A Proof of Theorem 1
Figure 10: Illustration of mining times of pacers and A-blocks after time $s$.
Under a confirmation rule of latency $t$, a target block is mined by time $s$. Within this proof, we count block arrivals from time $s$, i.e., the first A-block is the first A-block mined after $s$, and the first pacer is the first pacer mined after $s$. Let $X_{0}$ denote the time from $s$ to the mining time of the first A-block. For $i=1,2,\dots$, let $X_{i}$ denote the inter-arrival time between the $i$-th A-block and the $(i+1)$-st A-block. Then, $X_{0},X_{1},\dots$ are independent and identically distributed (i.i.d.) exponential random variables with mean $\frac{1}{a}$. Let $Y_{0}$ denote the time from $s$ to the mining time of the first pacer. For $i=1,2,\dots$, let $Y_{i}$ denote the inter-arrival time between the $i$-th pacer and the $(i+1)$-st pacer. It is easy to see that $Y_{1},Y_{2},\dots$ are i.i.d. random variables with mean $\frac{1}{h}+\Delta$, and $Y_{0}$ has mean less than $\frac{1}{h}+\Delta$. An illustration of these random variables is given in Fig. 10.
@@
$P(E_{j})$ $=P\Bigg{(}\sum_{i=0}^{j}X_{i}<\sum_{i=0}^{j}Y_{i},\sum_{i=0}^{j}X_{i}>t\Bigg{)}$ (56)
$\geq 1-P\Bigg{(}\sum_{i=0}^{j}(X_{i}-Y_{i})\geq 0\Bigg{)}-P\Bigg{(}\sum_{i=0}^{j}X_{i}\leq t\Bigg{)}$ (57)
$\geq 1-P\Bigg{(}\sum_{i=1}^{j}(X_{i}-Y_{i})\geq-n\Bigg{)}-P(X_{0}-Y_{0}\geq n)-P\Bigg{(}\sum_{i=0}^{j}X_{i}\leq t\Bigg{)}$ (58)
@@
$P\Bigg{(}\sum_{i=1}^{j}(X_{i}-Y_{i})\geq-n\Bigg{)}<\frac{\epsilon}{3}.$ (59)
@@ Appendix B Proof of Theorem 11
B-A Critical Events
Figure 11: Illustration of all possible state transitions over time.
Let a target block be given. As a convention, we refer to the genesis block as the 0-th balancing candidate. We describe several types of events as follows: For every $i\in\{0,1,\cdots\}$, let
@@ B-A Critical Events
If $U_{k}$ occurs, then at the end of the $k$-th candidacy, the candidate is not balanced, so the candidate becomes public and is strictly higher than all disagreeing blocks. As such, all honest mining power will extend chains that include the target transaction and are higher than the $k$-th candidate. This implies that the adversary must extend a shorter chain to catch up with the highest public honest block in order for the next candidate to be mined. In Appendix B-B, we show that the chance that no additional candidate will ever be mined is lower bounded as
$P(A_{k}|U_{k})\geq\delta$ (62)
@@ B-B Proof of (62)
$P\left(\cap_{t\geq 0}\{A_{r,r+t}\leq P_{\Delta,t}\}\right)=P\left(\cap_{t\geq 0}\{A_{0,t}\leq P_{0,t}\}\right)$ (63)
$=P\left(\max_{t\geq 0}\{A_{0,t}-P_{0,t}\}=0\right)$ (64)
$=P(M=0).$ (65)
@@
$P(A_{k}|U_{k})$ $\geq e^{-a\Delta}P(M=0)$ (66)
$=\delta.$ (67)
B-C Partially Markov Processes
@@
$P^{m}(\{0\,b_{1}\dots b_{k-1}0\})=P(\{0\,b_{1}\dots b_{k-1}0\})$ (71)
for $k=1,\dots,m-1$ and as
@@
for $k=m,m+1,\dots$. The outcome $00\,b_{1}\dots b_{k-1}$ 0 with $k\geq m$ implies that $C_{m}$ occurs. It is straightforward to verify that (71), (72) defines a pmf. Evidently, $P^{m}$ is partially Markov in the sense that it is consistent with $P$ up until stage $m-1$, after which $P^{m}$ becomes Markov with the same transition probabilities as $P^{0}$. In particular, conditioned on $C_{m}$, the balanced stages before stage $m$ and the balanced stages afterward are independent.
For every pair of natural numbers $l$ and $j$, we define the following pair of random variables as functions of the sample path $00\,b_{1}\dots b_{l-1}$ 0:
$L_{j}$ $=b_{1}+\dots+b_{\min(l,j)-1}$ (73)
$R_{j}$ $=b_{j}+\dots+b_{l-1},$ (74)
@@
In Appendices B-D and B-E, we show that for every $m=0,1,\dots$ and every $n=0,1,\dots$,
$P^{m+1}(X\leq n)\geq P^{m}(X\leq n)$ (76)
@@
$+P^{0}(R_{2}\geq n-1|B_{1})P^{0}(B_{1}|C_{1})$ (81)
$=$ $(1-\delta)(1-\epsilon)f_{n}+\epsilon f_{n-1}.$ (82)
@@
$f_{n}=\left(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\right)^{n}.$ (83)
@@
With confirmation by depth, there can be at most $k$ balanced heights before violation. In the remainder of this section, we are concerned with the number of balanced candidates up to the $k$-th candidacy, denoted as $X^{(k)}$. We first establish the following result:
$P(X^{(k)}\leq n)\geq P^{0}(X^{(k)}\leq n)$ (86)
where (86) can be proved as follows. Since we only consider the first $k$ candidates, which means only the first $k$ stages are relevant, we have
$P(X^{(k)}\leq n)=P^{k+1}(X^{(k)}\leq n).$
Furthermore, we can prove that for every $m=0,1,\dots,k$ and $n=0,1,\dots,k$,
$P^{m+1}(X^{(k)}\leq n)\geq P^{m}(X^{(k)}\leq n).$ (87)
For $X^{(k)}$, we define $L^{\prime}_{j}$ and $R^{\prime}_{j}$ as functions of the sample path $00\,b_{1}\dots b_{l-1}$ 0 for natural number $l$:
$L^{\prime}_{j}$ $=b_{1}+\dots+b_{\min(j,l)-1},\quad j=0,1,\dots,k,$ (88)
$R^{\prime}_{j}$ $=b_{j}+\dots+b_{\min\{k,l\}},\quad j=1,2,\dots,k+1,$ (89)
with boundary conditions $L^{\prime}_{0}=0$ and $R^{\prime}_{k+1}=0$. Thus, we can express $X^{(k)}$ as:
$X^{(k)}=L^{\prime}_{m}+1_{B_{m}}+R^{\prime}_{m+1},\quad\text{for}m\leq k.$ (90)
The remainder of the proof follows the same steps as in the proof of (76) in Appendix B-D, the details of which is omitted here.
Using (86), we can lower bound the cdf of $X^{(k)}$:
$P(X^{(k)}\leq n)$ $=P^{0}(X^{(k)}\leq n,X\geq n+1)+P^{0}(X\leq n)$ (91)
$=P^{0}(X^{(k)}\leq n,X\geq n+1,C_{k+1})+P^{0}(X\leq n)$ (92)
where (92) holds because $X^{(k)}\leq n$ and $X\geq n+1$ implies the occurrence of $C_{k+1}$. Since we already know $P^{0}(X\leq n)$ by (85), what left is to calculate $P^{0}(X^{(k)}\leq n,X\geq n+1,C_{k+1})$. By law of total probability, we have
$\begin{split}P^{0}(X^{(k)}\leq n,X\geq n+1,C_{k+1})=\sum_{i=0}^{n}P^{0}(X^{(k)}=i,C_{k+1})P^{0}(X-X^{(k)}\geq n+1-i|C_{k+1}).\end{split}$ (93)
@@
$=\sum_{i=0}^{n}\binom{k}{i}\epsilon^{i}((1-\epsilon)(1-\delta))^{k-i}(1-\delta)\Big{(}\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\Big{)}^{n-i+1}$ (94)
$=(1-\delta)\Big{(}\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\Big{)}^{n-k+1}\sum_{i=0}^{n}\binom{k}{i}\epsilon^{i}\Big{(}\frac{\epsilon(1-\epsilon)(1-\delta)}{\epsilon+\delta-\epsilon\delta}\Big{)}^{k-i}$ (95)
$=(1-\delta)\Big{(}\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\Big{)}^{n-k+1}\Big{(}\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\Big{)}^{k}\sum_{i=0}^{n}\binom{k}{i}(\epsilon+\delta-\epsilon\delta)^{i}((1-\epsilon)(1-\delta))^{k-i}$ (96)
$=(1-\delta)\Big{(}\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\Big{)}^{n+1}F_{b}(n;k,(\epsilon+\delta-\epsilon\delta)),$ (97)
@@ B-D Proof of (76)
$\begin{split}P^{m+1}(X\leq a)=P(L_{m}\leq a,C_{m}^{c})+\sum^{a}_{i=0}P(C_{m,i})P^{m+1}(X\leq a|C_{m,i}).\end{split}$ (107)
@@
$\begin{split}P^{m}&(X\leq a|C_{m,i})=P^{0}(B_{m},C_{m+1}|C_{m,i})x_{1}\\ &+P^{0}(U_{m},C_{m+1}|C_{m,i})x_{2}+P^{0}(A_{m}|C_{m,i}).\end{split}$ (115)
@@
$\lambda_{1}x_{1}+\lambda_{2}x_{2}+\lambda_{3}\leq\lambda_{1}^{\prime}x_{1}+\lambda_{2}^{\prime}x_{2}+\lambda_{3}^{\prime}.$ (118)
@@
$\lambda_{1}x_{1}+\lambda_{2}x_{2}+\lambda_{3}$ $=x_{2}+\lambda_{1}(x_{1}-x_{2})+\lambda_{3}(1-x_{2})$ (119)
$\leq x_{2}+\lambda_{1}^{\prime}(x_{1}-x_{2})+\lambda_{3}^{\prime}(1-x_{2})$ (120)
$=\lambda_{1}^{\prime}x_{1}+\lambda_{2}^{\prime}x_{2}+\lambda_{3}^{\prime}.$ (121)
@@
B-E Proof of (77)
By the definition of limit, proving (77) equals to prove $\forall\lambda\geq 0,a\in\{0,1,\dots\},\exists N_{0}$ such that $\forall N>N_{0}$, $|P^{N}(X\leq a)-P(X\leq a)|<\lambda$.
@@ B-E Proof of (77)
$P^{N}(X\leq a)$ $=P^{N}(C_{N}^{c})P^{N}(X\leq a|C_{N}^{c})+P^{N}(C_{N})P^{N}(X\leq a|C_{N}),$ (124)
$P(X\leq a)$ $=P(C_{N}^{c})P(X\leq a|C_{N}^{c})+P(C_{N})P(X\leq a|C_{N}).$ (125)
By definition of (71) and (72), $P(X\leq a|C_{N}^{c})=P^{N}(X\leq a|C_{N}^{c})$ and $P(C_{N}^{c})=P^{N}(C_{N}^{c})$. Let $p_{1}=P^{N}(X\leq a|C_{N})$ and $p_{2}=P(X\leq a|C_{N})$. We have,
$|P^{N}(X\leq a)-P(X\leq a)|$ $=P(C_{N})\,|p_{1}-p_{2}|$ (126)
$\leq(1-\delta)\left(1-\delta(1-\epsilon)\right)^{N-1}|p_{1}-p_{2}|$ (127)
$<\lambda,$ (128)
@@
$P^{0}(C_{N})$ $=P^{0}(C_{1}|U_{0})\prod_{i=1}^{N-1}P^{0}(C_{i+1}|C_{i})$ (132)
$=(1-\delta)\left(1-\delta(1-\epsilon)\right)^{N-1},$ (133)
@@ Appendix C Proof of Lemma 12
$X_{0}=F_{X}^{-1}(U+F_{X}(-a)-\lfloor U+F_{X}(-a)\rfloor)$ (142)
$Y_{0}=F_{Y}^{-1}(F_{Y}(a-1)-U-\lfloor F_{Y}(a-1)-U\rfloor)$ (143)
@@
$F_{X}(b)$ $\geq F_{X}\circ X_{0}(u)$ (147)
$=F_{X}\circ F^{-1}_{X}(u+F_{X}(-a))$ (148)
$>u+F_{X}(-a)\quad a.e.,$ (149)
@@
$F_{Y}(-b-1)$ $\geq F_{Y}\circ Y_{0}(u)$ (150)
$=F_{Y}\circ F^{-1}_{Y}(F_{Y}(a-1)-u)$ (151)
$>F_{Y}(a-1)-u\quad a.e.$ (152)